OpenStudy (anonymous):

solve the system by the substitution method 2x-5=y 3x-y=0

OpenStudy (anonymous):

x = -5, y = -15

OpenStudy (anonymous):

2x-y=5 3x-y=0 we'll multiply the bottom by -1 and add the two equations: 2x-y=5 -3x+y=0 -x=5 x=-5 plugging back into either equation y=-15